pleae help me in this
let abowl contain 10 chips of he same size and shape , one and only one of these chips is red , continue to draw chips from the bowl , one at atime and at random without replacement , until red chip is drawn ,
1. find the pdf. Of X , the number of trials needed to draw the red chip
2. compute p(X< or equal 4 )

Answer by stanbon(48568)

let abowl contain 10 chips of he same size and shape , one and only one of these chips is red , continue to draw chips from the bowl , one at atime and at random without replacement , until red chip is drawn ,
1. find the pdf. Of X , the number of trials needed to draw the red chip
P(1)=1/10
P(2)=(9/10)(1/9)=1/10
P(3)=(9/10)(8/9)(1/8)=1/10
...
P(10)=(9/10)(8/9)(7/8)...(1/2)(1/1)=1/10
---------
E(X)=(1/10)(1+2+3+...+9+10)=5.5
---------
2. compute p(X< or equal 4 )
=4/10
=================
Cheers,
Stan H.